A stone is thrown upwards from the ground with a velocity ‘u’
a) Obtain the maximum height attained by the stone?
b) Check the correctness of the equation obtained in (a) using the
method of dimensional analysis.
1
Expert's answer
2020-12-28T08:42:53-0500

"\text{a)the stone moves with equal slowness:}"

"V = V_0-gt"

"V_0=u"

"V=u-gt"

"\text{the maximum lifting height occurs when the stone begins to move down}"

"V=0"

"u-gt=0"

"t=\frac{u}{g}"

"\text{height of the stone above the ground:}"

"h = V_0t-\frac{gt^2}{2}=ut-\frac{gt^2}{2}"

"h_{max } \text{ when }t=\frac{u}{g}"

"h_{max}=u*\frac{u}{g}-\frac{g}{2}*\frac{u^2}{g^2}=\frac{u^2}{2g}"

"\text{b) dimensional analysis:}"

"h_{max}=\frac{u^2}{2g}"

"h_{max}-[m]"

"u-[m\/s]"

"g-[m\/s^2]"

"\frac{u^2}{2g}-[\frac{m^2/s^2}{m/s^2}]=[m]"

"\text{the dimensions of the left and right sides }"

"\text{of the formula coincide}"

Answer:"h_{max}=\frac{u^2}{2g}"